OpenStudy (anonymous):

Dan had feet of string. He cut off a piece that was feet long. How many feet of string did Dan have left? A. 1 feet B. 2 feet C. 2 feet D. 4 1 feet

jimthompson5910 (jim_thompson5910):

You're missing info. How much did he cut off and how long was the string originally?

OpenStudy (saifoo.khan):

incomplete.

OpenStudy (anonymous):

srry 3 1 over eight

OpenStudy (anonymous):

and one over eight

OpenStudy (anonymous):

he cut of a piece that was 1 and 3 over 4 feet long

jimthompson5910 (jim_thompson5910):

\[3\frac{1}{8}-1\frac{3}{4}\] \[3\frac{1}{8}-1\frac{6}{8}\] \[2\frac{9}{8}-1\frac{6}{8}\] \[1\frac{3}{8}\]

OpenStudy (anonymous):

thanks

OpenStudy (ar43r):

and the russian way: 3 1/8 - 1 3/4= 25/8 - 7/4= 25/8 - 14/8= 11/8 =1 3/8